# XAMBA: Enabling Efficient State Space Models on Resource-Constrained Neural Processing Units

### v3

@@ Abstract
State-Space Models (SSMs) have emerged as efficient alternatives to transformers for sequential data tasks, offering linear or near-linear scalability with sequence length, unlike transformers with quadratic-complexity attention. This makes SSMs ideal for long-sequence tasks in natural language processing (NLP), vision, and edge AI applications such as real-time transcription, translation, and contextual search. These applications demand lightweight, high-performance models for deployment on resource-constrained devices like laptops and PCs. While specialized accelerators have been proposed for emerging neural networks, designing new hardware is time-intensive, costly, and impractical for every model. Instead, optimizing models for existing neural processing units (NPUs) in AI PCs offers a scalable and efficient solution. Towards this end, we propose XAMBA, the first framework to enable and optimize SSMs on commercial off-the-shelf (COTS) state-of-the-art (SOTA) NPUs. Our approach follows a systematic three-step methodology: (1) enabling SSMs on NPUs, (2) optimizing performance to meet target Key Performance Indicator (KPI) requirements, and (3) trading accuracy for additional performance gains. After enabling SSMs on NPUs, XAMBA addresses key performance bottlenecks with two techniques: CumBA and ReduBA. These replace sequential CumSum and ReduceSum operations with matrix-based computations, significantly improving execution speed and memory efficiency. In addition, ActiBA further enhances performance by mapping computationally expensive activation functions (e.g., Swish, Softplus) to NPU hardware using piecewise linear approximations, reducing latency with minimal accuracy loss. Experimental evaluations on an Intel® Core™ Ultra Series 2 AI PC demonstrate that XAMBA achieves significant performance improvements, reducing execution latency by up to 2.6$\times$ compared to baseline implementation. Our code implementation is available at this link.
@@ 1 Introduction
SSMs, a classical framework for modeling dynamic systems through first-order differential equations, have gained prominence in machine learning for their efficiency in handling sequential data. Unlike transformers, which rely on quadratic-complexity attention mechanisms, SSMs achieve linear or near-linear scalability with sequence length by leveraging principles from convolutional and recurrent neural networks Gu et al. (2022). This makes SSMs highly suitable for long-sequence tasks like natural language processing, computer vision, and medicine, where they match transformers’ modeling capabilities with significantly lower computational overhead. Their efficiency is particularly critical for edge applications, such as personal assistants and real-time transcription, where SSMs enable transformative AI with reduced resource consumption and improved energy efficiency. Among SSM-based architectures, Mamba Gu & Dao (2024) and Mamba-2 Dao & Gu (2024) are the most prominent. Mamba introduces algorithmic innovations, such as selective scan, which efficiently processes sequential data by selectively updating only relevant states. Mamba-2 further refines this approach with the structured state-space duality (SSD) framework, which connects SSMs with attention mechanisms, enabling the reuse of transformer optimization techniques Patro & Agneeswaran (2024). These algorithmic advancements position SSMs as strong candidates for replacing transformers in resource-constrained environments. While specialized accelerators exist for emerging networks, designing a new hardware is costly and impractical for every model. Instead, XAMBA repurposes NPUs for SSMs, enabling efficient deployment on existing hardware. The efficiency of SSMs makes them ideal for deployment on NPUs, specialized accelerators optimized for data-parallel operations like matrix multiplication. Modern edge processors, such as AI PCs from Intel, Qualcomm, and AMD, integrate NPUs alongside CPUs and GPUs to support diverse AI workloads. NPUs, designed for high throughput and energy efficiency, consist of two key components: (1) Multiply-and-accumulate (MAC) Processing Units (MPUs) Raha et al. (2024), which handle parallelized matrix operations, and (2) Digital Signal Processors (DSPs), which execute sequential tasks like non-linear activations and cumulative summations. This architecture makes NPUs ideal for continuous, resource-intensive workloads such as personal assistants, real-time transcription, and contextual search. Prior work has demonstrated the benefits of deploying transformer-based models (e.g., Llama, Phi) on NPUs, achieving significant improvements in inference latency and energy efficiency Fei & Abdelfattah (2024). However, transformers’ quadratic complexity limits their scalability for long sequences. SSMs, with their linear complexity, offer a natural fit for NPUs, enabling even greater performance and energy efficiency. Despite this, deploying SSMs on NPUs presents unique challenges. Their sequential computations (e.g., CumSum, Swish, Softplus) and specialized operators misalign with NPUs’ data-parallel architecture, leading to inefficient DSP execution, increased memory traffic, and underutilized parallel units. While prior work has optimized recurrent models like LSTMs on NPUs Wang et al. (2018) and proposed specialized accelerators like MARCA for SSMs Li et al. (2024), no prior work addresses the challenges of deploying SSMs on COTS SOTA NPUs. As shown in Figure 1, SSMs face significant bottlenecks on NPUs. For Mamba, activation functions like Swish and Softplus OpenVINO Documentation (2024) dominate execution time due to inefficient DSP execution. In Mamba-2, CumSum and ReduceSum emerge as critical bottlenecks, exacerbated by poor memory reuse and increased off-chip traffic. These challenges underscore the need for specialized optimizations to fully leverage SSMs on NPUs. To address these challenges, we propose XAMBA, the first work to enable and optimize SSMs on COTS NPUs. XAMBA follows a systematic 3-step methodology: (1) enabling SSMs on NPUs, (2) optimizing performance to meet target KPI requirements, and (3) trading accuracy for additional performance gains.
@@
  • XAMBA achieves significant latency reductions for Mamba and Mamba-2 on an Intel® Core™ Ultra Series 2 NPU. CumBA reduces execution latency by 1.8$\times$, ReduBA by 1.1$\times$, and ActiBA by up to 2.6$\times$ compared to baseline unoptimized implementation.
@@ 4 Results
We evaluate XAMBA’s effectiveness in reducing inference latency for Mamba-based models. Figure 4(a) shows the average latency of a single-block Mamba-2 130M model on the Intel® Core™ Ultra Series 2 NPU. The CumBA technique reduces latency by 1.8×1.8$\times$1.8 × over the baseline, while ReduBA provides a $\times1$ reduction. Reduced inference latency directly translates to improved Tokens/second. Combining CumBA and ReduBA yields a 2.3×2.3$\times$2.3 × reduction, demonstrating the performance gains of XAMBA. Figure 4(b) presents the normalized inference latency breakdown for the Mamba-2 130M model, comparing the baseline with the optimized CumBA technique. In the baseline, CumSum operations contribute over 50% of the total latency. CumBA reduces this by transforming CumSum into matrix multiplication using a precomputed mask, achieving a 1.85×1.85$\times$1.85 × reduction in latency. Figure 4(c) shows the first inference latency breakdown for the Mamba 130M model with ActiBA optimizations. Using PLU for SoftPlus in a piecewise linear fashion reduces latency by 1.2×1.2$\times$1.2 ×. Further reductions (1.8×1.8$\times$1.8 ×) are achieved with SiLU, leading to a total 2.6×2.6$\times$2.6 × latency reduction with negligible quality loss when both SoftPlus and SiLU are mapped to PLU. We set a KPI target of 50 Tokens/s, following MobileLLM-125M Liu et al. (2024) (a comparable LLM), to ensure sufficient client-side responsiveness. With ActiBA optimizations, the Mamba-130M model’s decoding performance increases from 100 Tokens/s to 260 Tokens/s, surpassing the 50 Tokens/s KPI target. XAMBA’s optimizations, demonstrated on the 130M models, extend to larger models and inputs having similar bottlenecks, enabling comparable or greater speed-ups. While performance gains vary by model size and workload, the core principles hold. Ongoing work will further explore scalability and optimization for larger models.
@@ Acknowledgments
This work was supported in part by the Center for the Co-Design of Cognitive Systems (CoCoSYS) and the Center on Cognitive Multispectral Sensors (CogniSense), two research centers under the Joint University Microelectronics Program (JUMP) 2.0, a Semiconductor Research Corporation (SRC) initiative sponsored by DARPA.

### v4

@@ Abstract
State-Space Models (SSMs) have emerged as efficient alternatives to transformers for sequential data tasks, offering linear or near-linear scalability with sequence length, unlike transformers with quadratic-complexity attention. This makes SSMs ideal for long-sequence tasks in natural language processing (NLP), vision, and edge AI applications such as real-time transcription, translation, and contextual search. These applications demand lightweight, high-performance models for deployment on resource-constrained devices like laptops and PCs. While specialized accelerators have been proposed for emerging neural networks, designing new hardware is time-intensive, costly, and impractical for every model. Instead, optimizing models for existing neural processing units (NPUs) in AI PCs offers a scalable and efficient solution. Towards this end, we propose XAMBA, the first framework to enable and optimize SSMs on commercial off-the-shelf (COTS) state-of-the-art (SOTA) NPUs. Our approach follows a systematic three-step methodology: (1) enabling SSMs on NPUs, (2) optimizing performance to meet target Key Performance Indicator (KPI) requirements, and (3) trading accuracy for additional performance gains. After enabling SSMs on NPUs, XAMBA addresses key performance bottlenecks with two techniques: CumBA and ReduBA. These replace sequential CumSum and ReduceSum operations with matrix-based computations, significantly improving execution speed and memory efficiency. In addition, ActiBA further enhances performance by mapping computationally expensive activation functions (e.g., Swish, Softplus) to NPU hardware using piecewise linear approximations, reducing latency with minimal accuracy loss. Experimental evaluations on an Intel® Core™ Ultra Series 2 AI PC demonstrate that XAMBA achieves significant performance improvements, reducing execution latency by up to 4.8$\times$ compared to baseline implementation. Our code implementation is available at this link.
@@ 1 Introduction
SSMs, a classical framework for modeling dynamic systems through first-order differential equations, have gained prominence in machine learning for their efficiency in handling sequential data. Unlike transformers, which rely on quadratic-complexity attention mechanisms, SSMs achieve linear or near-linear scalability with sequence length by leveraging principles from convolutional and recurrent neural networks Gu et al. (2022). This makes SSMs highly suitable for long-sequence tasks Azizi et al. (2025) like natural language processing, computer vision, and medicine, where they match transformers’ modeling capabilities with significantly lower computational overhead. Their efficiency is particularly critical for edge applications, such as personal assistants and real-time transcription, where SSMs enable transformative AI with reduced resource consumption and improved energy efficiency. Among SSM-based architectures, Mamba Gu & Dao (2024) and Mamba-2 Dao & Gu (2024) are the most prominent. Mamba introduces algorithmic innovations, such as selective scan, which efficiently processes sequential data by selectively updating only relevant states. Mamba-2 further refines this approach with the structured state-space duality (SSD) framework, which connects SSMs with attention mechanisms, enabling the reuse of transformer optimization techniques Patro & Agneeswaran (2024). These algorithmic advancements position SSMs as strong candidates for replacing transformers in resource-constrained environments. While specialized accelerators exist for emerging networks, designing a new hardware is costly and impractical for every model. Instead, XAMBA repurposes NPUs for SSMs, enabling efficient deployment on existing hardware. The efficiency of SSMs makes them ideal for deployment on NPUs, specialized accelerators optimized for data-parallel operations like matrix multiplication. Modern edge processors, such as AI PCs from Intel, Qualcomm, and AMD, integrate NPUs alongside CPUs and GPUs to support diverse AI workloads. NPUs, designed for high throughput and energy efficiency, consist of two key components: (1) Multiply-and-accumulate (MAC) Processing Units (MPUs) Raha et al. (2024), which handle parallelized matrix operations, and (2) Digital Signal Processors (DSPs), which execute sequential tasks like non-linear activations and cumulative summations. This architecture makes NPUs ideal for continuous, resource-intensive workloads such as personal assistants, real-time transcription, and contextual search. Prior work has demonstrated the benefits of deploying transformer-based models (e.g., Llama, Phi) on NPUs, achieving significant improvements in inference latency and energy efficiency Fei & Abdelfattah (2024). However, transformers’ quadratic complexity limits their scalability for long sequences. SSMs, with their linear complexity, offer a natural fit for NPUs, enabling even greater performance and energy efficiency. Despite this, deploying SSMs on NPUs presents unique challenges. Their sequential computations (e.g., CumSum, Swish, Softplus) and specialized operators misalign with NPUs’ data-parallel architecture, leading to inefficient DSP execution, increased memory traffic, and underutilized parallel units. While prior work has optimized recurrent models like LSTMs on NPUs Wang et al. (2018) and proposed specialized accelerators like MARCA for SSMs Li et al. (2024), no prior work addresses the challenges of deploying SSMs on COTS SOTA NPUs. As shown in Figure 1, SSMs face significant bottlenecks on NPUs. For Mamba, activation functions like Swish and Softplus OpenVINO Documentation (2024) dominate execution time due to inefficient DSP execution. In Mamba-2, CumSum and ReduceSum emerge as critical bottlenecks, exacerbated by poor memory reuse and increased off-chip traffic. These challenges underscore the need for specialized optimizations to fully leverage SSMs on NPUs. To address these challenges, we propose XAMBA, the first work to enable and optimize SSMs on COTS NPUs. XAMBA follows a systematic 3-step methodology: (1) enabling SSMs on NPUs, (2) optimizing performance to meet target KPI requirements, and (3) trading accuracy for additional performance gains.
@@
  • XAMBA achieves significant latency reductions for Mamba and Mamba-2 on an Intel® Core™ Ultra Series 2 NPU. CumBA reduces execution latency by 2.7$\times$, ReduBA by 1.2$\times$, and ActiBA by up to 2.6$\times$ compared to baseline unoptimized implementation.
@@ 4 Results
Table 1: Quality of XAMBA model variants across benchmarks. Arrows indicate direction of better performance. Groups are separated by model size.
Model LAMBADA PPL $\downarrow$ HellaSwag ACC $\uparrow$ PIQA ACC $\uparrow$ ARC-E ACC $\uparrow$ ARC-C ACC $\uparrow$ OpenbookQA ACC $\uparrow$ Winogrande ACC $\uparrow$ AVG. ACC $\uparrow$
Mamba-130M $16.07$ $24.23$ $35.25$ $51.85$ $47.98$ $28.40$ $64.47$ $42.03$
Mamba-130M-PLU SiLU & Softplus $25.34$ $31.83$ $52.33$ $47.14$ $28.20$ $59.19$ $40.67$
Mamba2-130M $16.79$ $24.15$ $35.27$ $52.80$ $47.39$ $30.60$ $64.91$ $42.52$
Mamba2-130M-PLU SiLU & SoftPlus $16.78$ $24.15$ $35.27$ $52.80$ $47.43$ $30.60$ $64.91$ $42.53$
Mamba-370M $8.14$ $27.90$ $46.48$ $52.09$ $54.97$ $30.80$ $69.48$ $46.95$
Mamba-370M-PLU SiLU & Softplus $8.13$ $27.99$ $46.49$ $52.01$ $54.92$ $30.80$ $69.48$ $46.95$
Mamba2-370M $7.98$ $26.71$ $46.94$ $53.67$ $54.84$ $32.40$ $70.51$ $47.51$
Mamba2-370M-PLU SiLU & SoftPlus $7.98$ $26.71$ $46.94$ $53.83$ $54.84$ $32.40$ $70.51$ $47.54$
Mamba-790M $6.01$ $29.35$ $55.07$ $56.51$ $61.24$ $34.20$ $72.14$ $51.42$
Mamba-790M-PLU SiLU & Softplus $6.02$ $29.44$ $55.06$ $56.35$ $61.24$ $34.00$ $72.14$ $51.37$
Mamba2-780M $5.85$ $28.50$ $54.91$ $56.91$ $60.98$ $36.20$ $72.03$ $51.59$
Mamba2-780M-PLU SiLU & SoftPlus $5.85$ $28.50$ $54.92$ $56.91$ $61.03$ $36.20$ $72.03$ $51.60$
Mamba-1.4B $5.04$ $32.94$ $59.11$ $57.30$ $65.53$ $36.40$ $74.16$ $54.24$
Mamba-1.4B-PLU SiLU & Softplus $5.04$ $32.59$ $59.07$ $57.22$ $65.45$ $36.40$ $74.16$ $54.15$
Mmaba2-1.3B $5.02$ $33.11$ $59.94$ $58.25$ $64.14$ $37.80$ $73.23$ $54.41$
Mmaba2-1.3B-PLU SiLU & SoftPlus $5.02$ $33.19$ $59.93$ $58.33$ $64.14$ $37.80$ $73.23$ $54.44$
Mamba-2.8B $4.23$ $36.26$ $66.16$ $59.98$ $69.65$ $39.60$ $75.24$ $57.82$
Mamba-2.8B-PLU SiLU & Softplus $4.23$ $36.26$ $66.16$ $60.14$ $69.61$ $39.60$ $75.24$ $57.84$
Mmaba2-2.7B $4.09$ $36.26$ $66.58$ $61.72$ $69.57$ $38.80$ $76.39$ $58.22$
Mmaba2-2.7B-PLU SiLU & SoftPlus $4.09$ $36.26$ $66.60$ $61.80$ $69.61$ $38.80$ $76.39$ $58.24$
@@
We evaluate XAMBA’s effectiveness in reducing inference latency for Mamba-based models. Figure 4(a) shows the average latency of a single-block Mamba-2 130M model on the Intel® Core™ Ultra Series 2 NPU. The CumBA technique reduces latency by 2.7×2.7$\times$2.7 × over the baseline, while ReduBA provides a 1.2×1.2$\times$1.2 × reduction. Reduced inference latency directly translates to improved Tokens/second. Combining CumBA and ReduBA yields a 4.8×4.8$\times$4.8 × reduction, demonstrating the performance gains of XAMBA. Figure 4(b) presents the normalized inference latency breakdown for the Mamba-2 130M model, comparing the baseline with the optimized CumBA technique. In the baseline, CumSum operations contribute over 50% of the total latency. CumBA reduces this by transforming CumSum into matrix multiplication using a precomputed mask, achieving a 2.7×2.7$\times$2.7 × reduction in latency. Figure 4(c) shows the first inference latency breakdown for the Mamba 130M model with ActiBA optimizations. Using PLU for SoftPlus in a piecewise linear fashion reduces latency by 1.2×1.2$\times$1.2 ×. Further reductions (1.8×1.8$\times$1.8 ×) are achieved with SiLU, leading to a total 2.6×2.6$\times$2.6 × latency reduction with negligible quality loss when both SoftPlus and SiLU are mapped to PLU. Table 1 demonstrates that ActiBA’s hardware-friendly approximations introduce only negligible quality tradeoffs. For the original Mamba architecture, the largest accuracy drop occurs with the 130M model (42.03% $\to$ 40.67%, a 1.36% reduction), while larger models like the 1.4B version show remarkably stable performance (54.24% $\to$ 54.15%, just 0.09% difference). The Mamba2 variants exhibit even greater robustness to approximations—the 130M model’s accuracy remains virtually unchanged (42.52% $\to$ 42.53%), and interestingly, the 2.7B model shows a slight improvement (58.22% $\to$ 58.24%). These results collectively confirm that ActiBA’s optimizations successfully maintain model quality, with the maximum observed degradation being well under 1.5% even for the smallest model, while most variants experience changes of less than 0.1%. We set a KPI target of 50 Tokens/s, following MobileLLM-125M Liu et al. (2024) (a comparable LLM), to ensure sufficient client-side responsiveness. With ActiBA optimizations, the Mamba-130M model’s decoding performance increases from 100 Tokens/s to 260 Tokens/s, surpassing the 50 Tokens/s KPI target. XAMBA’s optimizations, demonstrated on the 130M models, extend to larger models and inputs having similar bottlenecks, enabling comparable or greater speed-ups. While performance gains vary by model size and workload, the core principles hold. Ongoing work will further explore scalability and optimization for larger models.
@@ Acknowledgments
This work was supported in part by the Center for the Co-Design of Cognitive Systems (CoCoSYS) and the Center on Cognitive Multispectral Sensors (CogniSense), two research centers under the Joint University Microelectronics Program (JUMP) 2.0, a Semiconductor Research Corporation (SRC) initiative sponsored by DARPA. The authors would like to thank Souvik Kundu (Intel Labs) and Zhifan Ye (Georgia Tech) for insightful discussions that contributed to this work.